ramishaya6:

A cyclist is competing in a race and decides to pass some fellow cyclists who are getting tired. The initial speed of the cyclist is 21km/h and he has 3.0X10^2 W of power. If the bicycle is able to convert 92% of the input energy into kinetic energy, how fast will the cyclist be travelling for 4.0 s? The combined mass of the cyclist and the bicycle is 78 Kg.

SemiDefinite:

E = Pt = 300 * 4 E_actual = 0.92 * E = 0.92 * 300 * 4 E_actual = change in KE = ΔΚΕ ΔΚΕ = 1/2 m (v^2 - u^2) = 1/2 * 78 * (v^2 - u^2) convert u=21km/h to m/s & solve